OpenStudy (anonymous):

Containing A(1, 3) and B(0, 2) Indicate the equation of the line meeting the given conditions. Please put the equation in standard form.

OpenStudy (campbell_st):

find the gradient of the line m = (3 - 2)/(1 - 0) the gradient is m = 1 use the point gradient formula y - 2 = 1( x -1) y = x +1 the standard form is x - y = -1

OpenStudy (anonymous):

Oh, that's how you do it. Thanks! :)